OpenStudy (anonymous):

Find the slope of the line that passes through the points (5, 8) and (9, 4). A. 3 B. 1 C. -1 D. four-thirds

OpenStudy (kc_kennylau):

slope=\(\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

A?

OpenStudy (kc_kennylau):

nope

OpenStudy (anonymous):

I got 2, but that's not an option.

OpenStudy (kc_kennylau):

just substitute the numbers into the formula first :)

OpenStudy (anonymous):

I still didn't get anything.

OpenStudy (kc_kennylau):

ok i'll help you substitute but you must cal the rest yourself

OpenStudy (kc_kennylau):

slope \(=\frac{4-8}{9-5}\)

OpenStudy (anonymous):

so c?

OpenStudy (kc_kennylau):

BINGO :D

OpenStudy (anonymous):

Thank you soooo much!. Could you help me with a few more if you have time please?

OpenStudy (kc_kennylau):

no problem and ofc :)

OpenStudy (anonymous):

Write an equation of a line in slope-intercept form that is parallel to x + 4y = 6 and passes through the point (-8, 5) A. y=−14x+3 B. y=−14x+6.75 C. y = x - 13 D. y = x + 13

OpenStudy (kc_kennylau):

Express the equation in point slope form first

OpenStudy (anonymous):

So would that be y=mx+b?

OpenStudy (anonymous):

Is it Y?

OpenStudy (anonymous):

Lol I mean B?

OpenStudy (kc_kennylau):

let me finish my lunch first sorry

OpenStudy (anonymous):

You're fine. No worries.

OpenStudy (kc_kennylau):

I think it's neither